Let V be the set of all vectors of the form (x1, x2, x3) in R

3

(a) x1 − 3x2 + 2x3 = 0.

(b) 3x1 − 2x2 + x3 = 0 and 4x1 + 5x2 = 0.

Find the dimension and basis for V.


1
Expert's answer
2021-03-31T11:10:12-0400

"v :={(x_1,x_2,x_3) \in R^3: 3x_1-3x_2+x_3=0}\\
basis\\
\begin{bmatrix}
 1 \\
 0\\
 0
\end{bmatrix},\begin{bmatrix}
 0 \\
 -3\\
 0
\end{bmatrix},\begin{bmatrix}
 0\\
 0\\
 2
\end{bmatrix}\\
dimension=3\\
b)
3x_1-2x_2+x_3=0\\
4x_1+5x_2=0\\
x_2=\frac{-4}{5}x_1\\
3x_1-2. \frac{-4}{5}x_1 +x_3=0\\
3x_1+\frac{8}{5}x_1 +x_3=0\\
\frac{23}{5}x_1 +x_3=0\\
x_3=\frac{-23}{5}\\
basis
\begin{bmatrix}
 1\\
 0\\
 0
\end{bmatrix},\begin{bmatrix}
 0\\
 \frac{-4}{5}\\
 0
\end{bmatrix},\begin{bmatrix}
 0\\
 0\\
 \frac{-23}{5}
\end{bmatrix}\\
dimension=3"